# SciLitLLM: How to Adapt LLMs for Scientific Literature Understanding

### v2

@@ 4.1.2. Baselines
  • CPT baselines: We compare SciLitLLM-base against leading open-source base models: ChatGLM (Zeng et al., 2024), Llama3 (AI@Meta, 2024), Llama2 (Touvron et al., 2023), Qwen2 (Yang et al., 2024), Qwen1.5 (Team, 2024), Mistral-7B (Jiang et al., 2023) and Mixtral-8x7B (Jiang et al., 2024).
@@
  • SFT baselines: We benchmark leading instruction LLMs including GPT-4o (OpenAI, 2023), GPT-3.5 (Brown et al., 2020), Llama3 (AI@Meta, 2024) and Qwen2 (Yang et al., 2024). We also report the performance of SciTulu-7B (Wadden et al., 2024), which is a fine-tuned Llama2-7B (Touvron et al., 2023) on SciRIFF.
4.2. Performance Overview (RQ1)
4.2.1. Base model performance
@@
The performance comparison of base models is shown in Table 2. SciLitLLM-base consistently outperforms other general base models across four scientific benchmarks. Specifically, compared with LLMs of less than 15 billion parameters, SciLitLLM-7B-Base shows an average accuracy improvement of 3.9% over Qwen2-7B. For LLMs with more than 50 billion parameters, SciLitLLM-72B-Base, with QLoRA training, outperforms all other LLMs (without quantization) as large as 110 billion parameters. The results demonstrate the effectiveness of CPT on high-quality scientific corpora, paving the way to a specialized LLM for scientific literature understanding.
4.2.2. Instruct model performance
@@
As shown in Table 3, SciLitLLM-7B-Instruct achieves the highest performance in 4 out of 5 domains on SciAssess, outperforming the second-best model by 3.6%. Notably, on SciRIFF, it surpasses baseline models by a substantial margin of 10.1%. Additionally, SciLitLLM-72B, trained using QLoRA, shows a 1.7% and 0.9% performance improvement over Qwen2-72B on SciAssess and SciRIFF, respectively.
@@
Detailed model performance on SciAssess is presented in Table 7, where SciLitLLM-7B and SciLitLLM-72B both lead in 12 and 13 out of 29 sub-tasks. Specifically, SciLitLLM-7B excels in tasks such as table extraction and molecule generation, likely benefiting from the comprehensive task coverage in our synthetic instruction dataset, SciLitIns. On SciRIFF, SciLitLLM-7B/SciLitLLM-72B ranks first in 8/6 out of 11 evaluations [^8] In SciRIFF, the Qasper and SciFact tasks have two different evaluation methods and thus two results..
4.3. Ablation Study (RQ2 & RQ3)
@@
We conducted ablation experiments on three key components in our pipeline: the CPT stage, the SFT data recipe, and the instruction quality filter, to demonstrate their effectiveness. It is important to note that all ablation experiments were performed on SciLitLLM-7B due to budget constraints.
4.3.1. Scientific knowledge injection via CPT (RQ2)
@@
Qwen2-7B-Instruct 48.8 48.7
Qwen2-7B-Base + SFT 48.1 51.6
Qwen2-7B-Base + CPT + SFT 52.4 59.4
Table 4. Ablation study of the CPT stage. The results demonstrate the effectiveness of the CPT stage in improving performance on scientific literature understanding.
We investigate the contribution of the CPT stage for SciLitLLM. We compare the three variants: (1) Qwen2-7B-Instruct: official instruct-model checkpoint; (2) Qwen2-7B-base + SFT: applying our SFT stage directly to Qwen2-7B-base without CPT; (2) Qwen2-7B-base + CPT + SFT: SciLitLLM-7B-Instruct.
As shown in Table 4, applying SFT alone to the Qwen2-7B-Base model does not lead to clear performance gains on SciAssess (-0.7%) and yields only a modest improvement on SciRIFF (+2.9%). In contrast, incorporating both CPT and SFT results in substantial performance enhancements: a 3.6% improvement on SciAssess and a 10.7% gain on SciRIFF. These results demonstrate that the CPT is crucial for effectively injecting scientific knowledge and significantly enhancing LLM performance on scientific literature understanding tasks.
4.3.2. Influence of SFT Data Recipes (RQ3)
@@
Infinity-Instruct 44.5 44.7
Infinity-Instruct + SciRIFF 42.2 53.9
Infinity-Instruct + SciRIFF + SciLitIns 52.4 59.4
Table 5. Ablation study of SFT data recipes.
We explore the influence of each ingredient in SFT data recipes. We incrementally add three datasets to the SFT training set: Infinity-Instruct, SciRIFF, and SciLitIns. As shown in Table 5, using only the Infinity-Instruct  results in the lowest performance on both SciAssess (44.5%) and SciRIFF (44.7%). This indicates that fine-tuning LLMs on general instructions alone is insufficient for scientific literature understanding, likely because Infinity-Instruct lacks specialized contents.
Adding SciRIFF to Infinity-Instruct improves performance on SciRIFF significantly but decreases performance on SciAssess. This discrepancy may be due to the disjoint coverage of scientific domains between SciRIFF and SciAssess. Finally, including SciLitIns along with Infinity-Instruct and SciRIFF boosts performance on both benchmarks, with SciAssess at 52.4% and SciRIFF at 59.4%. This demonstrates that including SciLitIns that covers less-represented scientific domains and tasks is beneficial for enhancing model performance in scientific literature understanding.
4.3.3. Influence of Instruction Quality Filter
@@
SciLitIns w/o filtering 51.1 56.2
SciLitIns w/ filtering 52.4 59.4
Table 6. Ablation study of SFT instruction quality filtering.
We conduct an ablation study to assess the impact of quality filter for synthetic instructions by varying whether the dataset SciLitIns  was filtered. As discussed in Section 3.2.2, this filter removes low-quality instructions evaluated from five key aspects. Table 6 shows that applying the filter significantly improved the performance of SciLitLLM-7B on SciAssess (+1.3%) and SciRIFF (+3.2%). This demonstrates that SFT quality filtering process effectively selects high-value educational instructions, thereby boosting the performance of SciLitLLMon scientific literature understanding.
5. Limitations
Despite the promising results achieved by SciLitLLM, there are several limitations that should be acknowledged:
  • Insufficient data volume. Compared with existing pre-training datasets (AI@Meta, 2024; Yang et al., 2024; Taylor et al., 2022), the amount of data used for CPT is not satisfying. Future work should consider incorporating a larger scientific corpus, potentially including scientific blogs or purely synthetic data, to enhance the model’s scientific knowledge base and overall performance.
  • Lack of reasoning enhancement. The current pipeline does not explore advanced reasoning techniques such as Chain-of-Thought (Wei et al., 2022b) or Tree-of-Thought (Yao et al., 2023) in the data construction or model inference stages. Investigating these methods could potentially improve the model’s inference capabilities and overall performance.
@@ 5. Limitations
6. Conclusion and Future Works
In this paper, we introduce SciLitLLM, a specialized model for scientific literature understanding. It is initialized with a general base model – Qwen2 (Yang et al., 2024), and trained through a sequential pipeline of continual pre-training (CPT) and supervised fine-tuning (SFT). For effective scientific knowledge injection during CPT, we propose model-based format and grammar correction method, along with text quality filtering measures. To ensure high-quality and diverse instructions during SFT, we devise instruction synthesis and quality control approaches. Our experiments on widely-used benchmarks demonstrate the effectiveness of this pipeline in adapting a general model to the field of scientific literature understanding. Specifically, SciLitLLM-7B achieves a 3.6% improvement on the SciAssess (Cai et al., 2024) and a 10.1% improvement on the SciRIFF (Wadden et al., 2024) compared to leading models with fewer than 10 billion parameters. SciLitLLM-72B, trained with QLoRA, also surpasses baseline open-source LLMs. We note that this pipeline could be easily adapted to other specialized domains, particularly those lacking adequate open-source corpora and high-quality instruction sets.
Our future work will focus on expanding the diversity and quality of the training data, as well as exploring more efficient methods for domain-specific knowledge injection and high-quality instruction generation. Moreover, we plan to expand our pipeline to include the RLHF (Ouyang et al., 2022) stage for better human preference alignment and enhanced safety.
@@ Appendix B CPT Quality Filter
We randomly select 50k samples from our 700k CPT data. These selected samples are then scored using the Llama3-70B model. The prompt utilized for this scoring process is as follows:
Prompt for CPT Data Quality Labelling Below is an extract from a textbook. Evaluate whether the text has a high educational value and could be useful in an educational setting for teaching from primary school to grade school levels using the additive 5-point scoring system described below. Points are accumulated based on the satisfaction of each criterion:
- Add 1 point if the extract provides some basic information relevant to educational topics, even if it includes some irrelevant or non-academic content like advertisements and promotional material. - Add another point if the extract addresses certain elements pertinent to education but does not align closely with educational standards. It might mix educational content with non-educational material, offering a superficial overview of potentially useful topics, or presenting information in a disorganized manner and incoherent writing style. - Award a third point if the extract is appropriate for educational use and introduces key concepts relevant to school curricula. It is coherent though it may not be comprehensive or could include some extraneous information. It may resemble an introductory section of a textbook or a basic tutorial that is suitable for learning but has notable limitations like treating concepts that are too complex for grade school students. - Grant a fourth point if the extract is highly relevant and beneficial for educational purposes for a level not higher than grade school, exhibiting a clear and consistent writing style. It could be similar to a chapter from a textbook or a tutorial, offering substantial educational content, including exercises and solutions, with minimal irrelevant information, and the concepts aren’t too advanced for grade school students. The content is coherent, focused, and valuable for structured learning. - Bestow a fifth point if the extract is outstanding in its educational value, and perfectly suited for teaching either at primary school or grade school. It follows detailed reasoning, the writing style is easy to follow, and offers profound and thorough insights into the subject matter, devoid of any non-educational or complex content.
After examining the extract: - Briefly justify your total score, up to 100 words. - Conclude with the score using the format: ”Educational score: ¡total points¿”
@@
Figure 6. Score distribution of the CPT Data
@@ Appendix C SFT Details
C.1. Instruction Generation Pipeline
@@
For each of the three scientific domains, we collect a set of high-impact research papers and construct a word frequency table. To generate a question in a given domain, we sample 20 keywords from the corresponding word table and insert them into the prompt for that question. To ensure fair representations of less frequent keywords, we use random sampling with a temperature setting of 3. We release our code, prompt templates, and word frequency tables. Below is an example of generating a table extraction question:
@@ C.2. Instruction Deduplication
C.3. Quality Assessment of Generated SFT Instructions
In section 3.2.2, we sample 10k instruction pairs from SciLitIns and evaluate them by Llama-3-70B using the below prompt.
@@ Appendix D Benchmark Details
  • SciRIFF (Wadden et al., 2024) evaluates essential scientific literature understanding capabilities, including information extraction, summarization, question answering, claim verification, and classification. Data points in SciRIFF are notable for their long input contexts and complicated structured outputs. The Qasper and SciFact tasks have two different evaluation methods and thus two results. We note that SciRIFF contains a separate training set used in the SFT stage in our study.
  • SciAssess (Cai et al., 2024) features an end-to-end benchmark of understanding PDF content. It includes 29 tasks from five scientific domains: fundamental science, alloy materials, biomedicine, drug discovery, and organic materials. SciAssess is used exclusively for testing in our evaluation.
@@ Appendix E Detailed Performance on SciAssess
Domain Task SciTulu-7B Mistral-7B Llama3-8B Qwen2-7B SciLitLLM-7B Llama3-70B Qwen2-72B SciLitLLM-72B GPT3.5 GPT4o
Fundamental Science Average 32.3 48.3 58.5 70.3 74.8 70.9 77.1 78.4 62.2 76.7
MMLU 35.5 52.1 59.4 64.0 65.3 76.2 78.7 78.5 64.3 84.2
CMMLU 27.5 31.8 46.1 79.2 87.8 65.5 86.6 89.6 44.9 78.7
Xz-Ch 34.6 51.7 64.7 71.7 75.4 73.4 74.0 75.5 73.2 73.4
Xz-En 31.5 57.6 63.6 66.2 70.7 68.4 69.1 69.9 66.5 70.3
[…]
AlloyQA 6.7 33.3 26.7 53.3 53.3 46.7 53.3 66.7 53.3 46.7
CompEx 9.0 9.0 8.1 10.1 7.2 34.7 19.3 19.5 24.8 50.5
TempEx 34.3 28.5 32.9 28.5 30.9 55.6 58.0 57.9 30.9 60.9
SampDiff 27.4 14.3 29.1 6.3 18.1 26.6 7.6 32.8 12.7 34.6
TreatSeq 42.2 54.9 67.6 65.7 68.6 60.8 75.5 68.6 38.2 67.6
[…]
BioQA 33.3 37.4 43.4 41.4 38.4 50.5 55.6 54.5 29.3 59.6
ChemER 68.3 93.2 84.3 92.1 90.5 86.1 90.9 91.4 93.3 90.3
DisER 80.8 82.2 80.9 87.2 88.4 79.8 81.7 80.9 90.5 81.1
CompDis 67.7 70.3 74.6 73.8 74.5 78.2 76.3 74.0 71.6 72.6
GeneFunc 70.9 79.9 88.8 92.9 89.6 87.1 85.5 81.3 88.8 96.4
GeneReg 85.9 92.8 92.1 97.1 95.9 95.9 95.9 95.7 94.2 93.7
[…]
AffEx 1.1 4.6 5.7 4.3 3.0 3.5 6.1 5.4 8.1 31.4
DrugQA 40.0 53.3 33.3 20.0 46.7 40.0 33.3 40.0 33.3 53.3
TagMol 7.3 0.0 10.2 1.5 20.1 23.0 13.0 27.1 0.6 7.8
MarkMol 28.4 15.9 18.0 34.4 32.8 53.3 31.7 52.4 48.8 63.8
MolDoc 44.0 46.0 50.0 56.0 56.0 48.0 50.0 58.0 44.0 54.0
ReactQA 25.3 25.3 32.6 28.4 26.3 50.5 36.8 32.6 34.7 37.9
ResTarg 31.9 66.2 74.3 77.0 47.7 72.3 77.3 77.3 47.4 55.7
[…]
ElecQA 26.0 20.0 41.0 30.0 28.0 33.0 49.0 33.0 26.0 68.0
OLEDEx 1.8 8.0 6.5 7.2 6.8 16.1 17.4 11.0 13.5 27.9
PolyQA 6.7 6.7 13.3 20.0 93.3 80.0 73.3 80.0 0.0 80.0
PolyCompQA 23.9 25.7 35.8 32.1 49.5 53.2 73.4 74.9 33.0 82.6
PolyPropEx 4.9 20.2 22.4 32.2 43.4 35.9 54.4 54.4 39.5 75.3
SolEx 26.2 31.8 34.0 35.7 32.9 36.2 42.3 38.5 35.8 45.9
ReactMechQA 27.3 31.8 18.2 40.9 18.2 36.4 59.1 48.2 22.7 59.1
Overall Average 33.2 40.6 45.0 48.8 52.4 55.7 57.8 59.5 45.5 63.4
Table 7. Detailed model performance on SciAssess tasks. SciLitLLM-7B shows significant improvement in the Fundamental Science and Organic Materials domains while maintaining comparable performance in other domains. Overall, SciLitLLM-7B achieves approximately 3.6% improvement over the second-best LLM.

### v3

@@ 4.1.2 Baselines
  • Base model baselines. We compare SciLitLLM-base against leading open-source base models: Llama-3.1 (AI@Meta, 2024), Qwen-2.5 (Team, 2024).
@@
  • Instruction model baselines. We benchmark leading instruction LLMs including GPT-[3.5, 4o] (Brown et al., 2020; OpenAI, 2023), Llama3.1-[8B, 70B] (AI@Meta, 2024) and Qwen2.5-[7B, 14B, 72B] (Yang et al., 2024; Team, 2024) and Mistral-7B (Jiang et al., 2023). We also report the performance of SciTulu-7B (Wadden et al., 2024), which is a fine-tuned Llama2-7B (Touvron et al., 2023) model on SciRIFF.
4.2 Performance Overview (Q1)
4.2.1 Base model performance
@@
The performance comparison of base models is shown in Table 2. SciLitLLM-base consistently outperforms other general base models across four scientific benchmarks. Specifically, compared with LLMs of less than 10 billion parameters, SciLitLLM-7B-Base shows an average accuracy improvement of 2.60% over Qwen2.5-7B. The results demonstrate that CPT on high-quality scientific corpora can significantly improve the LLMs’ ability to understand scientific contexts.
4.2.2 Instruct model performance
@@
As shown in Table 3, SciLitLLM-7B-Instruct achieves the highest performance in all 4 domains on SciAssess, outperforming the second-best model (i.e., Llama3.1-8B) by 4.0%. Notably, on SciRIFF, it surpasses baseline models by a substantial margin of 10.1%. Additionally, SciLitLLM-14B-Instruct shows a 4.6% and 7.5% performance improvement over Qwen2-14B-Instruct on SciAssess and SciRIFF, respectively. It even outperforms other open-source models with five times more parameters (i.e., Llama3.1-70B and Qwen2.5-72B) on both benchmarks. This demonstrates the effectiveness of our proposed pipeline in adapting a general-purpose LLM to a specialized model for scientific literature understanding.
@@
Detailed model performance on SciAssess is presented in Table 7, where SciLitLLM-7B lead in 15 out of 27 sub-tasks. Specifically, SciLitLLM excels in tasks such as table extraction and molecule generation, likely benefiting from the comprehensive task coverage in our synthetic instruction dataset – SciLitIns. On SciRIFF, SciLitLLM-7B and SciLitLLM-14B both rank first in 8 out of 11 evaluations [^8] In SciRIFF, the Qasper and SciFact tasks have two different evaluation methods and thus two results. For LLM-based evaluations in MuP and Qasper, we use GPT-4o instead of GPT-3.5, as was done in the original paper..
4.3 Ablation Study (Q2 & Q3)
@@
We conducted ablation experiments on three key components in our pipeline: the CPT stage, the SFT data recipe, and the instruction quality filter, to investigate their effectiveness. Note that all ablation experiments were performed on SciLitLLM-7B due to budget constraints.
4.3.1 Scientific knowledge injection via CPT (Q2)
@@
Qwen2.5-Instruct 46.5 50.3
Qwen2.5-Base+SFT 49.8 57.0
Qwen2.5-Base+CPT+SFT 51.7 60.6
Table 4: Ablation study of the CPT stage. The results shows that the CPT stage is essential to improve scientific literature understanding ability.
@@
Infinity-Instruct only 47.1 51.2
+ SciRIFF 46.8 56.7
+ SciRIFF + SciLitIns 51.7 60.6
Table 5: Ablation study of SFT data recipes. Our synthetic dataset SciLitIns effectively improves performance on both benchmarks.
@@
SciLitIns
w/o filtering 50.9 60.1
w/ filtering 51.7 60.6
Table 6: Ablation study of SFT instruction quality filtering. This shows that our proposed filter refines the quality of SciLitIns .
We investigate the contribution of the CPT stage for SciLitLLM. We compare the three variants: (1) Qwen2.5-7B-Instruct: official instruct-model checkpoint; (2) Qwen2.5-7B-base + SFT: applying our SFT stage directly to Qwen2.5-7B-base without CPT; (2) Qwen2.5-7B-base + CPT + SFT: SciLitLLM-7B-Instruct.
As shown in Table 6, applying SFT alone to the Qwen2.5-7B-Base model yields improvements on SciAssess (+3.3%) and SciRIFF (+6.7%). Further, incorporating CPT results in an extra 1.9% improvement on SciAssess and a 3.6% gain on SciRIFF. This shows the effectiveness of our CPT pipeline in curating high-quality CPT corpora and injecting scientific knowledge to LLMs.
4.3.2 SFT with SciLitIns (Q3)
We explore how fine-tuning with our synthetic instruction set, SciLitIns, impacts performance in scientific literature understanding. We incrementally add three datasets to the SFT training set: Infinity-Instruct, SciRIFF, and SciLitIns. As shown in Table 6, using only the general purpose instruction set (i.e., Infinity-Instruct) results in the lowest performance on both SciAssess (47.1%) and SciRIFF (51.2%). This indicates that fine-tuning LLMs on general instructions alone is insufficient for scientific literature understanding.
Adding the training set of SciRIFF to Infinity-Instruct improves performance on SciRIFF significantly (+5.5%) but decreases performance on SciAssess (-0.3%). This drop may be due to SciRIFF’s relatively narrower focus, primarily covering biomedicine and clinical medicine, while SciAssess spans broader fields such as biology, chemistry, and material science. This highlights the need for a diverse instruction set that covers diverse scientific domains. Finally, including SciLitIns boosts performance on both benchmarks, with SciAssess at 51.7% and SciRIFF at 60.6%. This demonstrates that SciLitIns, which is designed to contain instructions from less-represented scientific domains, is useful for improving model performance in scientific literature understanding.
4.3.3 Impact of Instruction filter
We conduct an ablation study to assess the impact of quality filter for synthetic instructions by varying whether the dataset SciLitIns  was filtered. As discussed in Section 3.2.2, this filter removes low-quality instructions evaluated from five key aspects. Table 6 shows that applying the filter improves the performance of SciLitLLM-7B on SciAssess (+0.8%) and SciRIFF (+0.5%). This demonstrates that SFT quality filtering process effectively excludes low-quality instructions, thereby boosting the performance of SciLitLLMon scientific literature understanding.
5 Limitations
Despite the promising results achieved by SciLitLLM, there are several limitations that should be acknowledged as follows:
  • Insufficient data volume. Compared with existing pre-training datasets (AI@Meta, 2024; Yang et al., 2024; Taylor et al., 2022), the amount of data used for CPT is comparatively small. Future work should consider incorporating a larger scientific corpus, potentially including scientific blogs or purely synthetic data.
  • Lack of reasoning enhancement. The current pipeline does not explore advanced reasoning techniques such as Chain-of-Thought (Wei et al., 2022) or Tree-of-Thought (Yao et al., 2023) in the data construction or model inference stages.
@@ 5 Limitations
6 Conclusion and Future Works
In this paper, we introduce SciLitLLM, a specialized model for scientific literature understanding. It is initialized with a general base model – Qwen2.5 (Yang et al., 2024), and trained through continual pre-training (CPT) and supervised fine-tuning (SFT). For effective scientific knowledge injection during CPT, we propose model-based format and grammar correction method, along with text quality filtering measures. To ensure high-quality and diverse instructions during SFT, we devise instruction synthesis and quality control approaches. Our experiments on widely-used benchmarks demonstrate the effectiveness of this pipeline in adapting a general model to the field of scientific literature understanding. Specifically, SciLitLLM-7B achieves a 4.0% improvement on the SciAssess (Cai et al., 2024) and a 10.1% improvement on the SciRIFF (Wadden et al., 2024) compared to leading models with fewer than 10 billion parameters. SciLitLLM-14B also surpasses leading open-source LLMs with around 70B parameters. We note that this pipeline could be easily adapted to other specialized domains, particularly those lacking adequate high-quality corpora and instruction sets.
Our future work will focus on expanding the diversity and quality of the training data, as well as exploring more efficient methods for domain-specific knowledge injection and high-quality instruction generation. Moreover, we plan to expand our pipeline to include the RLHF (Ouyang et al., 2022) stage for better human preference alignment and safety.
Ethics Statement
In developing SciLitLLM, we prioritized ethical considerations to ensure the responsible use of our models and methodologies. First, our research does not involve human subjects, and all data used for continual pre-training (CPT) are copyright-compliant. We employed a rigorous quality filtering process to minimize the risk of incorporating biased or misleading content. Nevertheless, we acknowledge that biases inherent in scientific literature, including historical underrepresentation of certain research domains, may propagate into the model’s outputs. We also adhere to all applicable legal and ethical research guidelines, such as respecting copyright policies during dataset construction and providing comprehensive model documentation. Our work is conducted with a commitment to research integrity, ensuring that our contributions remain beneficial to the scientific community while addressing ethical responsibilities associated with developing AI technologies.
Reproducibility Statement
We have made extensive efforts to ensure the reproducibility of our work. Our proposed SciLitLLM models are available anonymously. The corresponding training pipeline have been described in Sections 3. The source code for all modules shown in Figure 5 is included in the supplementary material. The training datasets for both CPT and SFT stages are shown in Table 1. We will release the generated instruction sets SciLitIns after administrative procedure.
@@ Appendix B CPT Quality Filter
We randomly select 50k samples from our CPT data. These selected samples are then scored using the Llama3-70B-Instruct model. The prompt utilized for this scoring process is as follows:
Prompt for CPT Data Quality Labelling Below is an extract from a textbook. Evaluate whether the text has a high educational value and could be useful in an educational setting for teaching from primary school to grade school levels using the additive 5-point scoring system described below. Points are accumulated based on the satisfaction of each criterion: - Add 1 point if the extract provides some basic information relevant to educational topics, even if it includes some irrelevant or non-academic content like advertisements and promotional material. - Add another point if the extract addresses certain elements pertinent to education but does not align closely with educational standards. It might mix educational content with non-educational material, offering a superficial overview of potentially useful topics, or presenting information in a disorganized manner and incoherent writing style. - Award a third point if the extract is appropriate for educational use and introduces key concepts relevant to school curricula. It is coherent though it may not be comprehensive or could include some extraneous information. It may resemble an introductory section of a textbook or a basic tutorial that is suitable for learning but has notable limitations like treating concepts that are too complex for grade school students. - Grant a fourth point if the extract is highly relevant and beneficial for educational purposes for a level not higher than grade school, exhibiting a clear and consistent writing style. It could be similar to a chapter from a textbook or a tutorial, offering substantial educational content, including exercises and solutions, with minimal irrelevant information, and the concepts aren’t too advanced for grade school students. The content is coherent, focused, and valuable for structured learning. - Bestow a fifth point if the extract is outstanding in its educational value, and perfectly suited for teaching either at primary school or grade school. It follows detailed reasoning, the writing style is easy to follow, and offers profound and thorough insights into the subject matter, devoid of any non-educational or complex content. After examining the extract: - Briefly justify your total score, up to 100 words. - Conclude with the score using the format: ”Educational score: ¡total points¿”
@@
Figure 6: Score distribution of the CPT Data
@@ Appendix C SFT Details
C.1 Instruction Generation Pipeline
@@
For each of the three scientific domains, we collect a set of high-impact research papers and construct a word frequency table. To generate a question in a given domain, we sample 20 keywords from the corresponding word table and insert them into the prompt for that question. To ensure fair representations of less frequent keywords, we use random sampling with a temperature setting of 3. We will release our code, prompt templates, and word frequency tables. Below is an example of generating a table extraction question:
@@ C.2 Instruction Deduplication
C.3 Quality Assessment of Generated SFT Instructions
In section 3.2.2, we sample 10k instruction pairs from SciLitIns and evaluate them by Llama-3-70B using the below prompt. Specifically, the quality is evaluated from five aspects: clarity, complexity, correctness, usefulness, and adaptability, assigning each instruction a score from 0 (lowest quality) to 5 (highest quality).
@@ C.3 Quality Assessment of Generated SFT Instructions
We show the quality statistics of synthetic instructions in Figure 7. Instructions with an average score of less than 4 are filtered out.
Figure 7: The quality of SciLitIns is evaluated from five aspects: clarity, complexity, correctness, usefulness, and adaptability (the higher the better). Instructions with an average score of less than 4 are filtered out.
@@ Appendix D Benchmark Details
  • SciRIFF (Wadden et al., 2024) evaluates essential scientific literature understanding capabilities, including information extraction, summarization, question answering, claim verification, and classification. Data points in SciRIFF are notable for their long input contexts and complicated structured outputs. The Qasper and SciFact tasks have two different evaluation methods and thus two results. For MuP and one evaluation method of Qasper, we use GPT-4o as language model evaluator, while the original paper uses GPT-3.5. We note that SciRIFF contains a separate training set used in the SFT stage in our study.
  • SciAssess (Cai et al., 2024) features an end-to-end benchmark of understanding PDF content. It includes 28 tasks from four scientific domains: biology, chemistry, materials and medicine. SciAssess is used exclusively for testing in our evaluation.
@@ Appendix E Detailed Performance on SciAssess
Domain Task SciTulu-7B Mistral-7B Llama3.1-8B Qwen2.5-7B SciLitLLM-7B Qwen2.5-14B Llama3-70B Qwen2.5-72B SciLitLLM-14B GPT3.5 GPT4o
Biology Average 45. 3 52.0 63.4 60.9 65.3 65.9 69.4 67.2 67.0 55.4 68.9
MmluPro 41.4 34.4 68.9 70.2 72.2 77.7 81.5 84.0 79.6 65.0 87.4
ChartQA 35.2 40.7 46.7 47.7 48.2 51.8 47.7 48.7 56.3 31.2 55.8
ChemRecog 54.9 56.0 73.7 66.7 84.3 80.1 83.6 76.4 77.8 64.9 79.5
CompRecog 41.6 72.4 69.3 66.8 72.2 74.5 76.8 71.2 71.1 63.6 73.3
DiseRecog 58.0 67.6 76.2 57.4 76.1 64.7 79.3 79.3 74.1 68.8 76.3
GeneFunc 40.6 40.7 45.8 56.6 38.9 46.4 47.4 43.8 43.2 39.1 41.0
Chemistry Average 19.0 34.0 46.1 47.9 55.4 60.0 62.5 63.7 63.4 37.3 68.0
MmluPro 13.8 17.9 40.8 46.2 52.6 63.7 67.6 72.3 63.2 30.3 74.5
ElecQA 19.6 41.7 53.8 63.3 70.4 75.9 75.5 78.5 78.4 30.5 85.5
OledExtr 0.9 22.7 13.7 36.3 30.6 42.0 56.3 49.9 37.4 28.0 43.8
ChartQA 33.3 46.7 86.7 46.7 66.7 80.0 80.0 80.0 80.0 66.7 86.7
PolyQA 25.0 56.7 59.1 76.9 88.0 90.9 85.2 91.4 91.3 33.0 93.8
PolyExtr 7.4 35.8 43.1 42.6 66.4 52.9 69.0 69.2 78.1 56.2 75.9
SolExtr 31.5 33.7 33.2 40.5 36.1 41.1 44.7 43.7 41.3 40.8 44.4
ReactQA 21.0 23.1 30.3 32.8 33.3 39.0 38.5 37.9 51.3 27.2 48.7
MechQA 18.2 27.3 54.5 45.5 54.5 54.5 45.5 50.0 50.0 22.7 59.1
Material Average 31.3 36.4 50.9 48.9 53.7 57.9 59.2 59.4 62.6 37.0 62.0
MatQA 36.9 43.0 59.7 56.7 58.6 66.5 73.8 71.9 70.3 52.1 76.8
ChartQA 33.3 66.7 66.7 73.3 46.7 60.0 46.7 53.3 53.3 40.0 46.7
CompExtr 8.0 8.9 9.9 13.9 40.6 37.1 45.7 43.0 42.9 18.9 46.2
TempQA 24.2 32.9 68.1 43.0 72.0 69.6 65.2 64.7 79.2 29.5 80.7
SampDiff 29.1 20.7 36.3 39.7 44.3 52.7 62.4 57.8 60.8 32.9 62.4
TreatSeq 56.4 46.5 64.9 66.8 59.9 61.4 61.4 65.8 68.8 48.5 59.4
Medicine Average 19.9 25.1 30.2 28.3 32.4 31.5 37.4 37.8 39.9 31.8 45.8
MmluPro 21.3 30.3 58.4 51.7 57.2 62.0 71.0 68.5 65.8 53.1 76.3
AffiExtr 2.7 3.7 3.7 2.6 5.6 4.0 4.7 7.1 3.9 5.5 10.1
ChartQA 33.3 40.0 26.7 40.0 33.3 33.3 40.0 33.3 53.3 33.3 46.7
TagMol 1.2 0.2 7.1 0.7 2.8 6.2 14.3 13.6 7.3 2.3 22.9
MarkMol 6.9 20.2 35.2 20.8 37.5 23.6 42.5 44.3 47.1 52.3 58.5
MolDoc 54.0 56.0 50.0 54.0 58.0 60.0 52.0 60.0 62.0 44.0 60.0
Table 7: Detailed model performance on SciAssess tasks.